OpenStudy (anonymous):

The maximum value of the expression 1/(sin^2 x + 3sinx cox + 5 cos^2 x)

OpenStudy (anonymous):

@satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73

OpenStudy (anonymous):

@amistre64 @amistre64 @amistre64 @amistre64 @amistre64 @amistre64 @amistre64 @amistre64

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

hi..@satellite73 plzz help

OpenStudy (amistre64):

the bottom looks like it could be quadratic withthe proper subs

OpenStudy (amistre64):

a^2 + 3ab + 5b^2

OpenStudy (amistre64):

or, since cos^2 = 1-sin^2, that might help reduce the clutter

OpenStudy (anonymous):

scratch that, it is wrong

OpenStudy (anonymous):

what amistre said same as \(4\cos^2(x)+3\cos(x)\sin(x)+1\) by replacing \(\sin^2(x)+\cos^2(x)=1\)

OpenStudy (anonymous):

apparently the minimum of the denominator is \(\frac{1}{2}\) so the max is \(2\) but i do not see why yet amistre?

OpenStudy (anonymous):

yup @satellite73 u r correct

OpenStudy (anonymous):

but hw did u get that

OpenStudy (anonymous):

no, wolfram is

OpenStudy (amistre64):

(sin^2 x + 3sinx cox + 5 cos^2 x)^(-1) -(sin^2 x + 3sinx cox + 5 cos^2 x)^(-2) * (2sinxcosx + 3cos^2x - 3 sin^2x - 5sinxcosx) -(sin^2 x + 3sinx cox + 5 cos^2 x)^(-2) * (3cos^2x - 3 sin^2x -3sinxcosx) = 0; when 3cos^2x - 3 sin^2x -3sinxcosx = 0; for extrema

OpenStudy (amistre64):

3cos^2x - 3 sin^2x -3sinxcosx cos^2x - sin^2x -sinxcosx = 0

OpenStudy (amistre64):

cos^2x - (1-cos^2x) - sin(2x)/2 = 0 cos^2x - 1 +cos^2x - sin(2x)/2 = 0 2cos^2x - sin(2x)/2 = 1 4cos^2x - sin(2x) = 2 etc ....

OpenStudy (anonymous):

unfortunately i seem to be getting the zero of the derivative at \(\tan^{-1}(\frac{1}{3})\)

OpenStudy (amistre64):

the wolf says we can simplify to: sin(2x) = 2cos(2x) tan(2x) = 2 2x = tan-1(2) x = tan-1(2)/2 abt:: .5536 rads, or aby 31 degrees

OpenStudy (amistre64):

i loathe trigging stuff lol

OpenStudy (anonymous):

i was just working with the denominator to find the min got the derivative via wolf as \(3\cos(2x)-4\sin(2x)\) which is a clue that the problem is cooked somehow, because we can write this as a single function of sine via \[3\cos(2x)-4\sin(2x)=-5\sin\left(2x-\tan^{-1}(\frac{3}{4})\right)\]

OpenStudy (anonymous):

but i am stuck from there this is zero when \[2x=\tan^{-1}(\frac{3}{4}) \]

OpenStudy (anonymous):

i think there must be some simpler way to tackle this